Accounting Help!Exercise 9-23 Budgeting Production and Direct-Material Purchases (LO 9-3, 9-6) Bodin Company budgets on an annual basis. The
Answer #1

1. Computation of number of units the firm would have to manufacture:

We know that:

Sales in units= Opening Inventory of Finished Goods + Production during the year - Closing Inventory of Finished Goods

468,000 = 97,000 + Production during the year - 76,000

Therefore, Production required during the year = 447,000

2. Computation of the amount of raw material needs to be purchased:

Note that as the opening and closing inventory of work in progress are same, it won't make an effect on raw material used or raw material to be purchased.

Total Raw material needed for the production during the year = Production of finished goods required * Units of raw material required for producing one unit of finished goods

Therefore, Total Raw material needed for the production during the year = 501,000 * 3 = 1,503,000 units

Raw material needed for the production = Opening Inventory + Raw Material Purchased - Closing Inventory

1,503,000 = 44,000 + Raw Material Purchased - 49,000

Therefore, Raw Material Purchased = 1,508,000 units.